# sample.bib

### v1

@@ Abstract
In this work, we introduce the concept of Active Representation Learning, a novel class of problems that intertwines exploration and representation learning within partially observable environments. We extend ideas from Active Simultaneous Localization and Mapping (active SLAM), and translate them to scientific discovery problems, exemplified by adaptive microscopy. We explore the need for a framework that derives exploration skills from representations that are in some sense actionable, aiming to enhance the efficiency and effectiveness of data collection and model building in the natural sciences.
@@ 1 Introduction
Representation learning is frequently used to uncover the latent structure in a data set and lies at the core of modern data science and machine learning. Deep learning-driven approaches have been widely applied in cases where data is abundant but challenging to interpret, for example in scientific data analysis. Frequently, such datasets cannot be labeled by human experts, because the nature of the scientific questions does not allow for clear labeling. For example, recent methods in the analysis of functional neuroimaging data rely on self-supervised learning techniques to uncover the geometric properties of the data set, often using methods like auto-encoding [Pandarinath2018-dp], or contrastive learning [Schneider2023-rn]. While these approaches can lead to remarkably insightful representations of complex data structures, they are mostly applied to scenarios where large, well-structured datasets are readily available for training general deep learning models.
However, in some data science problems, data has to be acquired by exploring a complex physical system whilst interacting with it through measurements. Consider, for example, active sensing-type problems, where data is collected through probing the input-output behavior of a black box process, including medical diagnosis, exploratory data analysis, active simultaneous localization and mapping (active SLAM), and, as we will explore throughout this work, adaptive light-sheet microscopy. Especially in light-sheet microscopy, we may find ourselves with the potential to collect, through a costly process, vast amounts of data of which only a small subset is truly informative about the questions and tasks at hand. Knowing how and where to collect new data can be challenging, but models of the underlying dynamical system and the measurement mechanism potentially help guiding the exploration decision towards informative regions of the space of possible measurements. However, generating a model requires integrating sequential decision making, in particular exploration with representation learning concepts. This raises the question: How can we integrate exploration and representation learning in a partially observable environment?
With the invention of Selective Plane Illumination Microscopy (SPIM; [huisken_2004]) the field of microscopy has entered the realm of real-time high data acquisition of biological processes, that opens the door for adaptive microscopy [Scherf2015-ji, Daetwyler2023-zu]. We see a strong analogy to problems in other areas such as robotics, in particular active SLAM (see e. g. [Placed2022-mf] for a recent review). This observation motivates a larger class of problems we call Active Representation Learning, sketched out in Figure 1.
@@ 2 Related Work
Exploration and Representation in Deep Reinforcement Learning.
Learning coherent state representations has been argued to be beneficial for deriving directed exploration strategies [yarats2021reinforcement]. In problems where the state space has a known structure, the representation learning problem can be simplified through inductive biases specific to the task and integrated with exploration strategies, as in visual active SLAM using Deep Reinforcement Learning [Chaplot_undated-lz, Chaplot2020-fn, Chaplot2020-vo]. While learning representations using world models [hafner2019learning, hafner2019dream], contrastive learning [laskin2020curl, stooke2021decoupling, yarats2021reinforcement], or data augmentation [laskin2020reinforcement] has demonstrated remarkable success in improving sample efficiency, they are usually solving a singular task. The emerging field of Unsupervised Reinforcement Learning (URL; [laskin2021urlb]) on the other hand, goes further by pre-training RL agents reward-free using intrinsic motivation objectives, derived from model-based curiosity metrics [Schmidhuber1991-tk, oudeyer2007intrinsic, pathak2017curiosity], or from information theoretic considerations [liu2021aps, laskin2022cic]. Other approaches involve Laplacian-like Representations [Erraqabi2021-xi, Klissarov2023-de, barreto2017successor], Skill-based Representations [eysenbach2018diversity, achiam2018variational], and combinations thereof [hansen2019fast, liu2021aps]. Exploration skills are usually obtained by deriving options [Sutton1999-gy] based on them [Machado2018-my, Klissarov2023-de]. Finally, integrating causal representation learning into the POMDP framework is another promising direction, where one is interested in disentangling controllable and uncontrollable latent factors of a partially observable environment [Sawada2018-ni, Lee_undated-dq, Sontakke2020-ts].
@@
The terms adaptive or smart microscopy are used for a wide range of solutions to common problems in microscopy and encompasses methods with the goal of maximizing the information content of the imaged volume. They aim for providing assistance throughout the imaging process. This includes optimizing the optical system during an imaging experiment  [Royer2016-dz, quirin_2016, McDole2018-ek] , managing and representing data [cheeseman_2018], applying processing techniques to improve image quality [guo_2020], detecting events of interest [mahecic_2022], [alvelid_2022], and incorporating prior knowledge [he_2020], [gritti_2022]. We here want to coin the term active microscopy where an underlying model of the whole dynamical system including the biological process under the microscope and the imaging mechanics is exploited to make active decisions that lead to high-quality observations of the sample.
@@ 3 Active Representation Learning
3.1 Motivation: From Active SLAM to Active Microscopy
The ability to uncover and use the latent structure of ones surroundings is an important aspect of intelligent and flexible behavior. Latent structure usually refers to states of a process that are not directly observable, but related in a regular way. Knowing this structure enables an autonomous controller to make educated guesses about the consequences of a course of action, even if that action has never been taken in this same context before. Guessing the consequence of an action is especially helpful, when interactions are costly, e. g. when they require energy as in robotic navigation, or when measurements are time-or storage-consuming. However, the structure that we hope to learn and exploit, e. g. a map of an environment, or a model of a biological process, is often hidden in the sense that it can only be observed through incomplete pieces of information, obtained through limited interactions.
@@ 3.1 Motivation: From Active SLAM to Active Microscopy
In microscopy, for example, the goal is often to uncover the 3D structure and dynamical processes of a biological specimens’ developmental trajectory. To this end, one tries to acquire a stack of images in the fastest and least invasive way possible. It might be tempting to blindly acquire as many planes as possible, however, imaging a biological process (e. g. a developing organism) is inherently subjected to the trade-off involving spatial and temporal resolution, signal-to-noise ratio and sample health. Additionally, absorption, scattering, and refraction lead to incomplete and noisy 2D image planes.
Modern light-sheet microscopes can quickly generate enormous amounts of data in the range of multiple terabytes, thus predicting which measurements are actually necessary would considerably simplify the data processing pipeline or even makes it feasible in the first place.
Considering these constraints, an active microscope should still be able to learn a useful representation of itself and its interaction with the biological process at hand. This involves disentangling the dynamical factors that are directly controllable (where and how to look), and those that are not (the biological process itself). Ideally, a useful model would include both in order to inform powerful exploration skills, e. g. how to effectively search the space of microscope parameters to obtain a high-quality image of an informative event.
3.2 Towards a Framework for Active Representation Learning
@@
The basic idea of Active Representation Learning (ARL) is that a data structure is in some sense embedded in a Partially Observable Markov Decision Process (POMDP; [Astrom1965-ym]). A POMDP is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R)$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ are the state, action, and observation spaces, and $T:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, $O:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$, and $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are the transition, observation and reward function respectively, see e. g. [littman_thesis]. Following active SLAM, we model ARL assuming a general belief-dependent utility, replacing the role of the reward function [Araya-Lopez_undated-qy].
Active SLAM introduces further specifications to the POMDP formalism in the form of a structured state-space. The state at time $t$ is usually assumed to be a combination of the robot’s pose $\mathbf{x}_{t}\in\mathcal{X}$ and a map $\mathbf{m}_{t}\in\mathcal{M}$ that the robot builds up over time while interacting with the POMDP. This results in a factored state space of the form $\mathcal{S}=\mathcal{X}\times\mathcal{M}$ [Placed2022-mf]. In standard active SLAM, the environment is usually taken to be static, i.e. there is no dynamics beyond the motion of the agent. This leads to simplified models, where the state change is directly caused by the agent’s intervention. Furthermore, the changes in $\mathbf{x}_{t}$ and $\mathbf{m}_{t}$ can be assumed to be independent, hence the agent can model its beliefs over its pose and the map independently.
@@
While, in active SLAM, $\mathbf{m}_{t}\in\mathcal{M}$ refers to a representation of 2D navigable space, we want it to refer instead to a more general notion of model that includes controllable and uncontrollable dynamic factors. Active microscopy factors as $\mathcal{S}=\mathcal{X}\times\mathcal{M}^{c}\times\mathcal{M}^{u}$, consisting of
@@
An additional difficulty to this problem, compared to standard active SLAM, is that while the sets that compose the total state space are factored, they may not be independent. For example, the quality of the image can and should influence our belief about the underlying biological process, which motivates the agent to find the best quality views.
In order to deal with the complexity of the problem, we propose the agent should have at least the following two components (Figure 1):
  1. 1.
    an actionable representation $\phi:\mathcal{H}\to\mathcal{Z}$, and
  2. 2.
    latent-conditional exploration skills $\pi_{z}:\mathcal{Z}\to\Delta(A)$.
The actionable representation is a mapping from the observational history to an encoding of the full state at time $t$ that respects the original factorization $\mathcal{Z}=\mathcal{Z}_{\mathcal{X}}\times\mathcal{Z}_{\mathcal{M}}^{c}\times\mathcal{Z}_{\mathcal{M}}^{u}$ and which can be used to derive outcomes of measurements of corresponding exploration skills. Factoring the representation and exploration policy in this way, may help simplify the problem and reveal analogies with representation learning. Furthermore, it makes sure we obtain a representation $\phi$ that aids interpretability and gives us a point of contact with the agent’s process model.
@@ 3.3 Challenges
Several issues make the formulation of a common decision framework difficult in this situation. On the one hand, a reward function in the traditional sense may not exist, e.g. it is not possible to formulate it as a function of state-action pairs, since it also depends on the internal belief state (see $\rho$-POMDP; [Araya-Lopez_undated-qy]). In such decision problems, the reward is then usually referred to as the agent’s utility, which can be composed of intrinsic and extrinsic motivations, as well as costs and constraints. Deriving suitable utility functions is an open problem that has received much attention in Unsupervised RL recently. How the latent state space should be represented by the agent and how uncertainty about it relates to the agent’s utility is another long-standing problem in active SLAM [Placed2022-mf], and in Representation Learning in Reinforcement Learning, see e.g. [Erraqabi2021-xi].
Beyond practical considerations, theoretical questions abound, involving the identifiability [Hyvarinen2023-gn] of the controllable and uncontrollable latent processes, and properties of the exploration behaviors that result from various utility functions. Finally, since we are usually dealing with physical dynamical systems, control theoretic issues can not be ignored, including controllability and reachability notions for the latent state-space, observability of the latent processes, and stability of exploration skills.
@@ 4 Conclusion
Active Representation Learning (ARL) presents a step forward in the intersection of exploration and representation learning within partially observable environments. By leveraging the concepts and techniques from active SLAM, we have illustrated how these ideas can be extended to a larger class of problems with a similar structure, exemplified by what we call active microscopy.
The proposed ARL framework emphasizes the importance of disentangling controllable and uncontrollable factors in the environment, thus enabling intelligent agents to make informed decisions about where and how to explore. We argue that this approach not only improves the quality of the data collected but also facilitates the creation of more robust and interpretable models of complex systems. Future research should focus on refining the theoretical foundations of ARL, addressing practical implementation issues, and exploring new applications across various domains. Ultimately, ARL holds the promise of transforming how we interact with and understand complex, dynamic systems, paving the way for breakthroughs in both artificial intelligence and scientific discovery.

### v2

@@ Abstract
In this work, we introduce Active Representation Learning, a class of problems that intertwines exploration and representation learning within partially observable environments. We extend ideas from Active Simultaneous Localization and Mapping (active SLAM), and translate them to scientific discovery problems, exemplified by adaptive microscopy. We explore the need for a framework that derives exploration skills from representations, aiming to enhance the efficiency and effectiveness of data collection and model building in the natural sciences.
@@ 1 Introduction
Representation learning is frequently used to uncover the latent structure in a data set and lies at the core of modern data science and machine learning. Deep learning-driven representation learning approaches have been widely applied in cases where data is abundant but challenging to interpret, for example in scientific data analysis. Frequently, such datasets cannot be labeled by human experts, because the nature of the scientific questions does not allow for clear labeling. For example, recent methods in the analysis of functional neuroimaging data rely on self-supervised learning techniques to uncover the geometric properties of the data set, often using methods like auto-encoding [Pandarinath2018-dp], or contrastive learning [Schneider2023-rn]. While these approaches can lead to remarkably insightful representations of complex data structures, they are mostly applied to scenarios where large, well-structured datasets are readily available for training general-purpose deep learning models.
However, in some data science problems, data has to be acquired by exploring a complex physical system whilst interacting with it through measurements. Consider, for example, active sensing-type problems, where data is collected through probing the input-output behavior of a black box process, including medical diagnosis, exploratory data analysis, active simultaneous localization and mapping (active SLAM), and adaptive light-sheet microscopy. Especially in light-sheet microscopy, we may find ourselves with the potential to collect, through a costly process, vast amounts of data of which only a small subset is truly informative about the questions and tasks at hand.
Knowing how and where to collect new data is challenging, but it can make representation learning more effective. In turn, having a representation of the underlying dynamical system and the measurement mechanism, potentially helps guiding data acquisition. However, such an active approach to representation learning requires integrating aspects of sequential decision making with representation learning concepts. This raises the question: How can we integrate representation learning and sequential decision making?
We observe that similar questions appear in a variety of problem domains, which motivates a larger class of problems we call Active Representation Learning, sketched out in Figure 1.
@@ 2 Related Work
ARL is inspired by the challenge of building an active microscope. With the invention of Selective Plane Illumination Microscopy (SPIM; [huisken_2004]) the field of microscopy has entered the realm of real-time high data acquisition of biological processes, that opens the door for adaptive microscopy [Scherf2015-ji, Daetwyler2023-zu]. The terms adaptive or smart microscopy are used for a wide range of solutions to common problems in microscopy and encompasses methods with the goal of maximizing the information content of the imaged volume. They aim for providing assistance throughout the imaging process, which includes optimizing the optical system during an imaging experiment  [Royer2016-dz, quirin_2016, McDole2018-ek] , managing and representing data [cheeseman_2018], applying processing techniques to improve image quality [guo_2020], detecting events of interest [mahecic_2022, alvelid_2022], and incorporating prior knowledge [he_2020, gritti_2022]. We here want to further coin the term active microscopy where an underlying model of the whole dynamical system including the biological process under the microscope and the imaging mechanics is exploited to make active decisions that lead to high-quality observations of the sample.
Representations in Online Learning
Representation learning and online learning can come together in different ways, two of which are particularly relevant to ARL. One problem formulation is to allow prior knowledge to improve sample efficiency, either in the form of given, pre-trained or shared representations [agarwal2023provablebenefitsrepresentationaltransfer, cella2023multitaskrepresentationlearningstochastic, yang2021impactrepresentationlearninglinear, zhang2022makinglinearmdpspractical] or through optimistic approaches to online learning [rakhlin2014onlinelearningpredictablesequences, flaspohler2021onlinelearningoptimismdelay]. While the theory behind these approaches is relevant to ARL, we want to learn to make decisions about how to best collect data for learning the representations in the first place. Another stream of work considers practical aspects of representation learning in continual settings, either with [hayes2019memoryefficientexperiencereplay, hayes2020remindneuralnetworkprevent] or without [rao2019continualunsupervisedrepresentationlearning, ren2022onlineunsupervisedlearningvisual] supervision. Continual unsupervised representation learning is an especially important aspects of ARL. However, it is not enough to address ARL, since it usually doesn’t consider the agent’s decision in acquiring new data for representation learning.
Representations in Deep Reinforcement Learning.
Learning coherent state representations has been argued to be beneficial for deriving directed exploration strategies [yarats2021reinforcement]. In problems where the state space has a known structure, the representation learning problem can be simplified through inductive biases specific to the task and integrated with exploration strategies, as in visual active SLAM using Deep Reinforcement Learning [Chaplot_undated-lz, Chaplot2020-fn, Chaplot2020-vo]. Learning representations using world models [hafner2019learning, hafner2019dream], contrastive learning [laskin2020curl, stooke2021decoupling, yarats2021reinforcement], or data augmentation [laskin2020reinforcement] has demonstrated remarkable success in improving sample efficiency in reinforcement learning. Perhaps strikingly, such representations have even been shown to drive behavior in the absence of a specific task or reward signal [laskin2021urlb]. ARL is motivated by exactly this kind of representation-driven decision making. While the works mentioned by [laskin2021urlb] demonstrate successful proofs of principle, they do not address the continual learning aspects mentioned earlier or partial observability, and their properties and limits are yet too poorly understood to be usable in real-world practical settings.
@@ 3 Active Representation Learning
3.1 Motivation
@@
Active Representation Learning is motivated by the concept of the active microscope. In developmental biology, the goal is often to uncover the 3D structure and dynamical processes of a biological specimens’ developmental trajectory. To this end, one tries to acquire a stack of images in the fastest and least invasive way possible. It might be tempting to blindly acquire as many planes as possible, however, imaging a biological process (e. g. a developing organism) is inherently subjected to the trade-off involving spatial and temporal resolution, signal-to-noise ratio and sample health. Additionally, absorption, scattering, and refraction lead to incomplete and noisy 2D image planes.
Modern light-sheet microscopes can quickly generate enormous amounts of data in the range of multiple terabytes, thus predicting which measurements are actually necessary would considerably simplify the data processing pipeline or even makes it feasible in the first place. Because we are seeking a concise representation of the biological process (see e. g.  [Kobayashi2022-jh]), the raw data is just a means to an end. An active microscope would be able to balance the imaging trade-offs, and learn a useful representation of the biological process while being as gentle with the specimen as possible [Scherf2015-ji]. This also involves learning a representation of the imaging process to inform the decision about how, where and when to look next for new data. This can be achieved by separating the dynamical factors that are directly controllable (where and how to look), and those that are not (the biological process itself), learning to control what can be controlled with minimal cost, and motivating the agent to curiously explore what can not.
Due to the cumulative nature of the costs incurred, and the temporal dependencies between the continual formation of representations and the agent’s decisions about how to operate the microscope, this is a problem at the intersection of sequential decision making and continual representation learning.
3.2 Relation to Other Decision Frameworks
If our only problem were to search for a good set of microscope parameters to obtain high-quality views of the biological specimen, one would rightfully consider this a bandit task. It is already known from previous works (see Sec.  2), that external knowledge about the data generating process (e.g. in the form of representations) can be shown to improve sample complexity of decision algorithms. However, what sets the ARL problem apart is that state representations are not only unknown, learning them is the primary goal, and deriving a suitable learning objective is part of the problem.
During data acquisition, the potentially incomplete representation must be used as a basis for informing the next action to take, similar to how the incompleteness of the map in active SLAM is used to inform where to go next. This circular dependency of representation and decision making is the basis of ARL and its framing as a generalization of active SLAM. In general, ARL is difficult to frame as a bandit, since the agent-environment system has a complex state.
3.3 From Active SLAM to Active Microscopy
@@
The basic idea of Active Representation Learning (ARL) is that a data structure is in some sense embedded in a Partially Observable Markov Decision Process (POMDP; [Astrom1965-ym], see Appendix A.1). ARL is inspired by active SLAM, where we interpret the map as an example of a data structure embedded in a POMDP. Following SLAM, we model ARL assuming a general belief-dependent utility instead of an external reward [Araya-Lopez_undated-qy].
@@
Active SLAM introduces further specifications to the POMDP formalism in the form of a structured state-space. The state at time $t$ is usually assumed to be a combination of the robot’s pose $\mathbf{x}_{t}\in\mathcal{X}$ and its internal representation (a map) $\mathbf{m}_{t}\in\mathcal{M}$ of the external state space. This results in a factored state space of the form $\mathcal{S}=\mathcal{X}\times\mathcal{M}$ [Placed2022-mf]. While, in active SLAM, $\mathbf{m}_{t}\in\mathcal{M}$ refers to a representation of 2D navigable space, we want it to refer instead to a more general notion of model that includes controllable and uncontrollable dynamic factors. Active microscopy, for example, factors as $\mathcal{S}=\mathcal{X}\times\mathcal{M}^{c}\times\mathcal{M}^{u}$, consisting of
@@
A setup like this is not only reminiscent of active SLAM, but also of reinforcement learning-based visual attention [minut2001]. We can easily think of the microscope setup as a robotic environment in the context of RL. The controllable parts of the microscope act like a body, whereas the imaged volume is the outside environment. Our agent has to learn how to influence its observations, and how to do so in order to perceive the most interesting aspects of the environment’s uncontrollable parts.
@@ 3.4 Challenges
An immediate difficulty is that there is no specific reward function. While we can choose from a wide variety of intrinsic motivation objectives from the literature, most are not stationary functions of state-action pairs, but depend on some internal belief state (see $\rho$-POMDP; [Araya-Lopez_undated-qy]). Deriving suitable utility functions is an open problem that has received much attention in Unsupervised RL recently, and is also a known problem in SLAM. How the latent state space should be represented by the agent and how uncertainty about it relates to the agent’s utility is another long-standing problem in active SLAM [Placed2022-mf], and also in Reinforcement Learning, see e.g. [Erraqabi2021-xi]. The non-stationarity and partial observability of the resulting decision problem, as well as the continuity of representation learning present significant challenges for practical implementation. Beyond practical considerations, theoretical questions abound, involving the identifiability [Hyvarinen2023-gn] of the controllable and uncontrollable latent processes, and properties of the exploration behaviors that result from various utility functions. Finally, since we are usually dealing with physical dynamical systems, control theoretic issues can not be ignored, including controllability and reachability notions for the latent state-space, observability of the latent processes, and stability of exploration skills.
@@ 4 Conclusion
Active Representation Learning (ARL) presents a step forward in the intersection of decision making and representation learning within partially observable environments. By leveraging the concepts and techniques from active SLAM, we have illustrated how these ideas can be extended to a larger class of problems with a similar structure, exemplified by what we call active microscopy. The proposed ARL framework emphasizes the importance of disentangling controllable and uncontrollable factors in the environment, thus enabling intelligent agents to make informed decisions about where and how to explore. We argue that this approach not only improves the quality of the data collected but also facilitates the creation of more robust and interpretable models of complex systems. Future research should focus on refining the theoretical foundations of ARL, addressing practical implementation issues, and exploring new applications across various domains. Ultimately, ARL holds the promise of transforming how we interact with and understand complex, dynamic systems, paving the way for breakthroughs in both artificial intelligence and scientific discovery.
@@ Acknowledgments
The authors would like to thank Simon Hirländer for insightful comments on the practical issues in ARL, and the anonymous reviewer for valuable feedback on the connections of ARL with existing methods. N. M. and N.S. are supported by BMBF (Federal Ministry of Education and Research) through ACONITE (01IS22065) and the Center for Scalable Data Analytics and Artificial Intelligence (ScaDS.AI.) Leipzig. N.M. is also supported by the Max Planck IMPRS CoNI Doctoral Program. J.H. is supported by the Alexander von Humboldt Foundation (Alexander von Humboldt Professorship; J.H.) and the German Research Foundation (Germany’s Excellence Strategy EXC 2067/1-390729940; J.H.). G.M. is supported by the MWK (Niedersächsisches Ministerium für Wissenschaft und Kultur, 6707040) and is member of the Hertha Sponer College of the Cluster of Excellence Multiscale Bioimaging (MBExC; EXC 2067/1- 390729940), University of Goettingen, Germany. Figure 1 was created using BioRender.com.
@@
Appendix A Extended Background
A.1 Background on POMDPs and Active SLAM
A Partially Observable Markov Decision Process (POMDP; [Astrom1965-ym]) is defined by the tuple $(\mathcal{S},\mathcal{A},\mathcal{O},T,O,R)$, where $\mathcal{S}$, $\mathcal{A}$, and $\mathcal{O}$ are the state, action, and observation spaces, and $T:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$, $O:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{O})$, and $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ are the transition, observation and reward function respectively, see e. g. [littman_thesis].
The goal is to design a policy $\pi:\mathcal{H}\to\mathcal{A}$ where $\mathcal{H}$ is the set of histories of observations and actions, i.e. $\pi(a_{t}|o_{t},a_{t-1},o_{t-1},...)$, to maximize the expected reward:
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\middle|\pi\right]$
with $\gamma\in[0,1)$ being a discount factor. As opposed to the fully observable case, the decision maker cannot reliably determine their own true state, $s$, based on the current observation alone, and a Markovian optimal policy does not exist in general. However, the agent can summarize the history by computing a belief state $b\in\mathcal{B}=\Delta(S)$. Ideally, it would compute the Bayesian update
$b(s^{\prime}|a,o)\propto O(o|s^{\prime},a)\int_{s\in S}T(s^{\prime}|s,a)b(s),$
however, this is usually intractable to compute, and e.g. variational approximations are used in practice.
Since the Bayesian belief state serves as a sufficient statistic of the history, optimal policies for the original POMDP can be determined by solving an equivalent continuous-space MDP, known as a belief MDP $(\mathcal{B},\mathcal{A},T_{b},r_{b})$, where the new transition $T_{b}$ and reward functions $r_{b}$ are defined over ×$\mathcal{B}\times\mathcal{A}$. Using belief MDPs, several theoretical results about MDPs can be extended to POMDPs, such as the existence of optimal deterministic policies.
In this continuous MDP, the goal is to maximize the cumulative reward by identifying a policy that uses the current belief state as input. Formally, we seek a policy $\pi^{*}$ that satisfies
$\pi^{*}=\arg\max_{\pi}\mathbb{E}\left[\sum_{t=0}^{\infty}\gamma^{t}r_{b}(b_{t},a_{t})\right],$
where $r_{b}(b_{t},a_{t})=\int_{s\in\mathcal{S}}b(s_{t})r(s_{t},a_{t})$.
While POMDPs are very general, they can be difficult to map onto active sensing-type problems. Active SLAM introduces further specifications to the POMDP formalism in the form of belief-dependent rewards [Araya-Lopez_undated-qy], and a structured state-space. The state at time $t$ is usually assumed to be a combination of the robot’s pose $x_{t}\in\mathcal{X}$ and a map $m_{t}\in\mathcal{M}$ that the robot builds up over time while interacting with the POMDP, resulting in a factored state space of the form $\mathcal{S}=\mathcal{X}\times\mathcal{M}$.